# Understanding Nonlinear Collaboration between Human and AI Agents: A Co-design Framework for Creative Design

### 14 Jan 2024 v1

@@ 5. Study II: Evaluation of OptiMuse
The purpose of Study II is to evaluate the usability of OptiMuse and validate the effectiveness of the novel human-AI collaboration framework. We examine users’ design modification processes when performing specific tasks on given graphical designs while utilizing two tools.
Figure 6. User interactions in OptiMuse follow these steps: (1) users review slides on the preview window; (2) users input commands and engages in rule-based conversation with OptiMuse; (3) OptiMuse generates alternatives as selective choices upon reaching convergence with users; (4) users remix alternatives through rule-based conversation; (5) users are satisfied with the alternatives and choose one option after several rounds of steps 1 to 4 (Notes: A1-10 refer to the AI agent actions defined in Sec. 4.1).
Figure 7. An overview of the scenario of Study II.
@@ 5.1.1. Choice of Modification Tasks and Design Materials
We choose Wizard of Oz (WoZ) experiments for Study  II. The reasons are threefold. First, Copolit was not available since it had not been released to the public when we conducted Study II. Second, our target is to assess the two workflows of co-design collaboration. WoZ experiment allows us to manage the agent’s output as expected under specific conditions. Third, in WoZ environments, we can control the precise timing of each task. It is important to introduce supplementary constraints through scenarios in WoZ experiments (Norman1991wizard, ; dahlback1993wizard, ; green2004wizard, ; Riek2012wizardscenario, ). To do so, we analyzed the graphical designs, corresponding requirements, and requirement-operational strategies chain collected from Study I. First, we categorized the modification requirements based on levels of ambiguity and feasibility. Then, we selected two specific tasks in graphical design: modifying the color scheme and adjusting the text layout. The reasons are as below. (i) These tasks possess a certain degree of ambiguity, requiring exploration and investigation. (ii) They limit the scope of modification possibilities, which makes it feasible to predict the requirement analysis process. (iii) They frequently existed in Study I, suggesting their natural occurrence in the co-design practice. To avoid unequal familiarity with designs, users were required to use different co-design tools complecting the same task on different designs. Upon selecting two tasks, we carefully prepared two visually appealing graphical designs that were suitable for both tasks.
@@ 5.1.2. Wizard
To ensure a valid WoZ simulation, we adhered to the guidance provided by Fraser and Gilbert (Norman1991wizard, ) in selecting, regulating, and training the wizard. First, a qualified co-design assistant in our study requires a strong design background. Thus, we enlisted an experienced design professional with extensive experience as a design leader in the university students’ union for 2 years. This individual possesses a strong background in effectively communicating design requirements with supervisors and assisting fellow colleagues in accomplishing modification tasks. Second, to make the output from the wizard resemble that of a computer as far as possible regarding timing and consistency, we have established specific guidelines for the assistant to minimize randomness during the communication and recommendation process. Furthermore, the assistant was instructed to adopt certain behavior patterns commonly associated with AI agents. This was done to create an authentic interaction experience for participants, enhancing the perception that they were truly interacting with AI agents. Third, to ensure consistency in understanding, communication skills, and response time, we intentionally recruited only one wizard for our user study. In Appendix B, we provide additional details of guidelines for the wizard.
@@ 5.2.2. Baseline
We conducted a comparison between OptiMuse and Copilot (Fig. 8), which is a prototype lacking a dialogue function and generates a single option at a time. Much like other widely used AI-powered design tools, including the publicly released demos of Microsoft PowerPoint Copilot(copilot, ). Copilot operates by executing commands.
@@ 5.2.3. Procedure
Group ID tool design Tasks tool design Tasks
1 OptiMuse design 1 Task 1, 2 Copilot design 2 Task 1, 2
2 OptiMuse design 2 Task 1, 2 Copilot design 1 Task 1, 2
3 Copilot design 1 Task 1, 2 OptiMuse design 2 Task 1, 2
4 Copilot design 2 Task 1, 2 OptiMuse design 1 Task 1, 2
Upon completing the four tasks, the participants were requested to fill out two System Usability Scale (SUS) questionnaires (SUS, ) to assess the usability of two AI agents (OptiMuse and Copilot). At the end of the experiment, we conducted a follow-up interview and asked each participant the following questions: (i) Reviewing this experiment, describe the challenges you encountered while using Copilot and OptiMuse. (ii) What are the advantages and disadvantages of using OptiMuse compared to Copilot in co-design tasks? Please clarify the differences first. What are their performances in handling requirements ambiguity, conflicts, and changes? How do selective choices make a difference? How does Q&A make a difference? (iii) What specific assistance would you expect from AI agents like Copilot and OptiMuse in co-design? Are there any tasks or aspects that you prefer to handle independently, considering that they might exceed the capabilities of AI assistants? If so, what are the reasons behind your preference? (iv) Which collaboration mode is more similar to human behavior? What are the differences and pros and cons between the two collaboration modes and human collaboration? The study lasted 90 minutes. No participant expressed doubt regarding the presence of a wizard behind AI agents.
@@ 5.3. Results
Figure 9. Dialogue turns and average time in finishing one task for each participant in using OptiMuse and Copilot.
Figure 10. System Usability Scale (SUS) score of each participant shows a divergent attitude toward OptiMuse.
@@ 5.3.1. Quantitative findings
The average SUS score for OptiMuse is 70.21 (Good), while for Copilot, it is 64.17 (Poor). This indicates an improvement in the usability of our nonlinear human-AI collaboration as compared to the conventional linear approach (Fig. 10). We adhere to the instructions provided in the reference(results, ) for presenting our quantitative findings in time and dialogue turns (Fig. 9). Dialogue turns were computed by counting how many commands each participant sent to the AI agent. $P_{4}$ is an outlier who happened to be quite fond of the graph provided by OptiMuse, so the process concluded quickly. After removing $P_{4}$, we obtained the following results: (i) Participants spent an average time of 1292 seconds and 747 seconds using OptiMuse and Copilot to complete one task, respectively. A Wilcoxon Signed-rank test shows that there is a significant effect of tools (W = 2, Z = -2.76, p ¡ 0.05, r = -0.9). (ii) The average turns of OptiMuse and Copilot were 10 and 8 rounds. A Wilcoxon Signed-rank test shows that there is a significant effect of tools (W = 10, Z = -2.04, p ¡ 0.05, r = -0.6). (iii) $P_{1,5,7,8}$ gave up on tasks when using Copilot as they lost confidence in AI’s ability to understand their design intentions and found it annoying in wording after an average of 1031 (365, 1803, 659, 1297) seconds and 8 (3, 10, 5, 15) turns. In conclusion, time, dialogue turn, and failure rate show that participants had greater tolerance toward OptiMuse in comparison to Copilot (Fig. 9).
5.3.2. People’s divergent attitudes towards OptiMuse and their interaction patterns
@@
Both SUS results 10 and interviews show a divergence in attitudes towards OptiMuse. While some participants were inspired by OptiMuse’s textual feedback and selective choices, others exhibited signs of anxiety and reluctance to engage in conversation.
@@
Positive feedback: fostering creativity and emulating human interaction. Individuals who praised OptiMuse were pleasantly surprised by the way it reshaped their initial design objectives. As $P_{10}$ mentioned: “ I was a little surprised when OptiMuse pointed out the problem of my requirements and offered alternatives. But it seems okay to give it a shot.” They became receptive to OptiMuse’s feedback, maintaining an open-minded perspective. They came to trust OptiMuse due to its logical and actionable textual feedback, as well as some gratifying selective choices. Despite the presence of conflicts between texts and visual output or misunderstandings, participants were enthusiastic about engaging with OptiMuse and allowing it to undergo trial and error. “ Indeed it appeared that OptiMuse did not do exactly what he said, but at the very least, such feedback provided me with a fundamental idea of what I could expect. ” We noticed some interesting interaction patterns in these users:
  1. (1)
@@
    Asking questions: We noticed users frequently inquiring about OptiMuse’s opinion in the dialogue. When reluctant to articulate potential operations, they inquired: “ What steps can I take in this scenario? Any suggestions?” When seeking feedback and validation, their query was: “ How do you view my proposals?” When faced with the complexity of discerning disparities among various options, they posed the question: “ What contrasts exist among these alternatives?”
  2. (2)
@@
    Remixing alternative choices: We found participants got inspiration and new design ideas by remixing their requirements. Actions of remixing, as a commonly acknowledged approach to fostering creativity, were also spotted in our participants. There were mainly two means to achieve remixing among several choices. First, remixing was observed through textual means. Our research yielded findings consistent with those of Kulkarni et al. (kulkarni2023word, ). Textual content lends itself to a more straightforward form of remixing. During active discussions of potential operations, users amalgamate various ideas by rephrasing them into a new description. A simple action of copy worked well. Second, remixing involves adding commands to multiple choices, aligning with designers’ inclination to draw insights from various sources. When they found satisfaction in specific aspects of different choices, they endeavored to identify those elements and merge them into a more improved solution. However, despite describing the custom of engaging in such interactions based on past design practices and emphasizing their importance due to prior experiences with co-design AI tools, $P_{5}$ didn’t actually carry out these interactions. This was because he believed it wasn’t possible, largely because popular AI tools didn’t support such actions.
@@
Negative feedback: too much AI and time waste. Some users had little tolerance for misunderstandings and unsatisfactory results. “I simply don’t require these interactions. I’m not concerned about how it comprehends my requirements. My priority is to see results directly and as quickly as possible.” $P_{4}$ added, “In fact, since communicating with AI is so complex, why not handle it myself?” Usually, they knew exactly how to operate, and they found it frustrating to translate their ideas into detailed descriptions. $P_{1}$ expressed: “OptiMuse was always trying to figure out what I meant by adding details or making descriptions more operated. I was kind of dizzy seeing the sentences became longer and longer…” Additionally, $P_{9}$ reported a lack of professional vocabulary, which affected users’ willingness to communicate about graphic design in the dialogue. Some interesting interaction patterns were noticed in these users:
  1. (1)
@@
    Avoiding chats: These users strongly opposed engaging in conversations with OptiMuse and devised various methods to circumvent dialogue. Some of them added “Let’s Have a try!” after their commands to prompt OptiMuse to generate outcomes without engaging in conversations during the study. Once they discovered this trick, they employed it consistently. Another tactic involved typing ahead and anticipating responses. Some users never changed what they typed based on the actual replies. These users were actually iterating ideas by their own writing because they paid no attention to OptiMuse’s feedback.
  2. (2)
@@ 5.3.3. Characteristics of Requirements in Design Practices
Although people’s attitudes varied a lot towards OptiMuse, we observed consistency in the characteristics of requirements among users using the same system. Participants agreed that the characteristics of requirements were unconsciously influenced by the differences between the two frameworks.
@@
Copilot: Strive for clarity, embrace change, and prevent conflicts. As we mentioned in Section 3, ambiguity is almost unavoidable in requirement descriptions. However, we observed how people grapple with the challenge of being specific when using Copilot, as evidenced in their dialogues. Additionally, during interviews, they self-reported the difficulty of being as detailed as possible to ensure Copilot’s understanding. $P_{1}$ mentioned, “I was uncertain about how much detail I should provide for Copilot to ensure accurate understanding.” The requirements underwent slight changes to achieve a suitable degree or shift the focus to another aspect of the design. The background was initially instructed to be “darker,” followed by “a little bit darker,” then “still not dark enough,” and finally, “well, try to make it a little lighter.” Similar iterations of descriptions were frequently employed in dialogues between humans and Copilot. Conflicts of the requirement could not be well fixed in Copilot. First, conflicts between asynchronous requirements like “darker and brighter” were regarded as AI’s lack of ability to produce appropriate outcomes. Second, conflicts between the user and AI, namely, AI didn’t give outcomes addressing the user’s commands, were regarded as annoying mistakes. “I aimed for white titles, but for some reason, it consistently incorporated a gradient, which I found displeasing.“, $P_{10}$ stated. Conflicts within the design itself, the most tricky ones, were usually avoided in discussions. $P_{3}$ explained, “When I requested Copilot to justify both sides of all body text, the result was quite poor. I usually adjust the text box width when aligning the text. But I didn’t bother explaining this to the AI.“
@@
OptiMuse: Ambiguity sparks alternative ideas, while changes and conflicts are the crucibles of creativity. There was a deliberate use of intentionally ambiguous descriptions observed in the interactions between participants and OptiMuse, as indicated by their dialogues. Furthermore, participants expressed their willingness to take OptiMuse’s opinions into account, reducing their concern about the precision of descriptions. $P_{3}$ said,“I was curious about OptiMuse’s take on this design, so I provided a vague description of my ideas and awaited its response.” OptiMuse was designed to highlight changes in requirements and conflicts, which is often where participants reported gaining novel design ideas. “I was surprised when OptiMuse questioned my commands, but its suggestions appeared reasonable. So I decided to give it a try,” $P_{8}$ added, “At least one of the selections turned out to be satisfactory. To be honest, I have never thought of doing so.” OptiMuse drove users to change requirements. In OptiMuse, users’ textual descriptions allow rapid exploration of the design space (kulkarni2023word, ). OptiMuse’s suggestions (transitioning from dark mode to white mode, switching theme colors, rearranging each text module based on different rules, etc) increased designers’ willingness to make bold changes. Users were prompted to decide among the changes, which encouraged reflection on each option. Users took greater tolerance towards conflicts with OptiMuse. Attitudes towards asynchronous conflicts were comparable in Copilot and OptiMuse. Nonetheless, such conflicts occurred less frequently in OptiMuse, as users were more focused on exploring new options rather than fine-tuning details. Conflicts between AI and users’ commands still happened in visual outcomes, but rather than simply dismissing them as mistakes, people attempted to contemplate the merits of selective choices. “It felt like having an opinionated colleague. You know, negotiating and presenting options. So, I suppose there must be reasons behind whatever OptiMuse produces.”, $P_{5}$ stated. Some participants were also receptive to OptiMuse’s alternative suggestions during the dialogue.
Table 3. Four expected roles of AI agents and their corresponding tasks.
[…]
@@
Roles
Input
Interaction
Output
Executor
Design
Confirm, Revise
Optimized Design
Optimizer
Design + Requirement
Confirm, Revise
Optimized Design
Colleague
Design + Requirement
Dialogue, Select, Remix
Selective Choices
Expert
Design + Requirement
Dialogue, Select, Remix
Selective Choices

### 20 Mar 2024 v3

@@ 5. Study II: Evaluation of OptiMuse
The purpose of Study II is to evaluate the usability of OptiMuse and validate the effectiveness of the novel human-AI collaboration framework. We examine users’ design modification processes when performing specific tasks on given graphical designs.
@@ 5.1.1. Choice of Modification Tasks and Design Materials
We use Wizard of Oz (WoZ) experiments for Study  II. The reasons are threefold. First, Copolit was not available since it had not been released to the public when we conducted Study II. Second, our target is to assess the two workflows of co-design collaboration. WoZ experiment allows us to manage the agent’s output as expected under specific conditions. Third, in WoZ environments, we can control the precise timing of each task. It is important to introduce supplementary constraints through scenarios in WoZ experiments (Fraser and Gilbert, 1991; Dahlback et al., 1993; Green et al., 2004; Riek, 2012). To do so, we analyzed the graphical designs, corresponding requirements, and requirement-operational strategies chain collected from Study I. First, we categorized the modification tasks (requirements) based on levels of ambiguity and feasibility. Then, we selected two specific tasks in graphical design: modifying the color scheme and adjusting the text layout because (i) These tasks possess a certain degree of ambiguity, requiring exploration and investigation; (ii) They limit the scope of modification possibilities, which makes it feasible to predict the requirement analysis process; (iii) They frequently existed in Study I, suggesting their natural occurrence in the co-design practice. To avoid unequal familiarity with designs, users were required to use different co-design tools to complete the same task on different designs. Upon selecting two tasks, we carefully prepared two visually appealing graphical designs that were suitable for both tasks.
@@ 5.1.2. Wizard
To ensure a valid WoZ simulation, we adhered to the guidance provided by Fraser and Gilbert (Fraser and Gilbert, 1991) in selecting, regulating, and training the wizard. A qualified co-design assistant in study II requires a strong design background. Thus, we enlisted an experienced design professional with extensive experience as a design leader in the university students’ union for 2 years. This individual possesses a strong background in effectively communicating design requirements with supervisors and assisting fellow colleagues in accomplishing modification tasks. Second, to ensure A1-10 while making the output from the wizard resemble that of a computer as far as possible regarding timing and consistency, we established specific guidelines for the assistant to minimize randomness during the communication and recommendation process. Furthermore, the assistant was instructed to adopt certain behavior patterns commonly associated with AI agents. This was done to create an authentic interaction experience for participants, enhancing the perception that they were interacting with an AI agent. Third, to ensure consistency in understanding, communication skills, and response time, we recruited only one wizard user II Additional details of guidelines for the wizard are provided in Appendix B.
Figure 5. An overview of the scenario of Study II.
5.2. Study Design: A Wizard-of-Oz Study
@@ 5.2.3. Baseline
We conducted a comparison between OptiMuse (Fig. 7) and Copilot (Fig. 6), which is a prototype lacking the dialogue function and generating a single outcome each time. Much like other widely used AI-powered design tools, including the publicly released demos of Microsoft PowerPoint Copilot(Microsoft, 2023), our Copilot operates by executing commands.
Figure 7. Overview of the OptiMuse : (a) The preview window shows the current outcome; (b) Users can communicate with OptiMuse through dialogues; (c) Copilot produces multiple visual outcomes when users type “Let’s have a try!”.
@@ 5.2.4. Procedure
Upon completing the four tasks, the participants were requested to fill out two System Usability Scale (SUS) questionnaires (Trend, 2024) to assess the usability of two tools (OptiMuse and Copilot). At the end of the experiment, we conducted a follow-up interview and asked each participant the following questions: (1) Reviewing this experiment, describe the challenges you encountered while using Copilot and OptiMuse. (2) What are the advantages and disadvantages of using OptiMuse compared to Copilot in co-design tasks? Please clarify the differences first. What are their performances in handling requirements ambiguity, conflicts, and changes? How do selective choices make a difference? How does Q&A make a difference? (3) What specific assistance would you expect from AI agents like Copilot and OptiMuse in co-design? Are there any tasks or aspects that you prefer to handle independently, considering that they might exceed the capabilities of AI assistants? If so, what are the reasons behind your preference? (4) Which collaboration mode is more similar to human behavior? What are the differences and pros and cons between the two collaboration modes and human-human collaboration? The study lasted 90 minutes. No participant expressed doubt regarding the presence of a wizard behind AI agents.
@@
Group tool design Task tool design Task
1 OptiMuse 1 1, 2 Copilot 2 1, 2
2 OptiMuse 2 1, 2 Copilot 1 1, 2
3 Copilot 1 1, 2 OptiMuse 2 1, 2
4 Copilot 2 1, 2 OptiMuse 1 1, 2
@@ 5.3.1. Quantitative findings
The average SUS score for OptiMuse is 70.21 (Good), while for Copilot, it is 64.17 (Poor). This indicates an improvement in the usability of our nonlinear human-AI collaboration as compared to the conventional linear approach.
@@
We adhere to the instructions provided in the reference(Yatani, 2014) for presenting our quantitative findings in time and dialogue turns. Dialogue turns were computed by counting how many commands each participant sent to the AI agent. $P_{4}$ is an outlier who happened to be quite fond of the design provided by OptiMuse, so the process concluded quickly. After removing $P_{4}$, we obtained the following results: (i) Participants spent an average time of 1292 seconds and 747 seconds using OptiMuse and Copilot to complete one task, respectively. A Wilcoxon Signed-rank test shows that there is a significant effect of tools (W = 2, Z = -2.76, p ¡ 0.05, r = -0.9). (ii) The average turns of OptiMuse and Copilot were 10 and 8 rounds. A Wilcoxon Signed-rank test shows that there is a significant effect of tools (W = 10, Z = -2.04, p ¡ 0.05, r = -0.6). $P_{1,5,7,8}$ gave up on tasks when using Copilot as they lost confidence in AI’s ability to understand their design intentions and found it annoying in wording after an average of 1031 (365, 1803, 659, 1297) seconds and 8 (3, 10, 5, 15) turns. In conclusion, time, dialogue turn, and failure rate show that participants had greater tolerance toward OptiMuse in comparison to Copilot (Fig. 8).
Figure 8. (a) Dialogue turns and (b) average time in finishing one task for each participant using OptiMuse and Copilot. (c) The System Usability Scale (SUS) score of each participant shows a divergent attitude toward OptiMuse (Notes: Underlined participants think OptiMuse is better than Copilot in terms of SUS score).
5.3.2. Divergent attitudes towards OptiMuse and the interaction patterns
@@
Both SUS results (Fig. 8) and interviews show divergent attitudes towards OptiMuse. While some participants were inspired by OptiMuse’s textual feedback and selective choices, others were anxious and reluctant to engage in conversation.
@@
Positive feedback: fostering creativity and emulating human interaction. Individuals who praised OptiMuse were pleasantly surprised by the way it reshaped their initial design objectives. As $P_{10}$ mentioned: “ I was a little surprised when OptiMuse pointed out the problem of my requirements and offered alternatives. But it seems okay to give it a shot.” They became receptive to OptiMuse’s feedback, maintaining an open-minded perspective. They came to trust OptiMuse due to its logical and actionable feedback, as well as some gratifying selective choices. Despite the presence of conflicts between texts and visual output or misunderstandings, participants were enthusiastic about engaging with OptiMuse and allowed it to undergo trial and error. “ Indeed it appeared that OptiMuse did not do exactly what he said, but at the very least, the textual feedback provided me with a fundamental idea of what I could expect.  ” We noticed some interesting interaction patterns in these users:
@@
  • Asking questions: Users frequently inquired about OptiMuse’s opinion in the dialogue. When reluctant to articulate potential operations, they asked: “ What steps can I take in this scenario? Any suggestions?” When seeking feedback and validation, their asked: “ How do you view my proposals?” When faced with the complexity of discerning disparities among various options, they asked: “ What contrasts exist among these alternatives?”
@@
  • Remixing alternative choices: Participants got inspiration and new design ideas by remixing design choices. Actions of remixing, as a commonly acknowledged approach to fostering creativity, were also frequently spotted in our participants. There were mainly two means to achieve remixing among several choices. First, remixing was observed through textual means. Our research yielded findings consistent with those of Kulkarni et al. (Kulkarni et al., 2023). Textual content lends itself to a more straightforward form of remixing. During active discussions of potential operations, users amalgamate various ideas by rephrasing them into a new description. A simple action of copy worked well. Second, remixing involves adding commands to multiple choices, aligning with designers’ inclination to draw insights from various sources. When they found satisfaction in specific aspects of different choices, they endeavored to identify those elements and merge them into a more improved solution. However, despite describing the custom of engaging in such interactions based on past design practices and emphasizing their importance due to prior experiences with co-design AI tools, $P_{5}$ didn’t actually carry out remixing of multiple selections. This was because he believed it wasn’t possible (despite our instruction), largely because popular AI tools didn’t support such actions.
@@
Negative feedback: too much AI and time waste. Some users had little tolerance for misunderstandings and unsatisfactory results. “I simply don’t require these interactions. I’m not concerned about how it comprehends my requirements. My priority is to see results directly and as quickly as possible.” $P_{4}$ added, “In fact, since communicating with AI is so complex, why not handle it myself?” Usually, they knew exactly how to modify the design, and found it frustrating to translate their ideas into detailed descriptions. $P_{1}$ expressed: “OptiMuse was always trying to figure out what I meant by adding details or making descriptions more operated. I was kind of dizzy seeing the sentences became longer and longer…” Additionally, $P_{9}$ reported a lack of professional vocabulary, which affected users’ willingness to communicate about graphic design in the dialogue. Some interesting interaction patterns were noticed in these users:
@@
  • Avoiding chats: These users strongly opposed engaging in conversations with OptiMuse and devised various methods to circumvent dialogue. Some of them added “Let’s Have a try!” after their commands to prompt OptiMuse to generate outcomes without engaging in conversations during the study. Once they discovered this trick, they employed it consistently. Another tactic involved typing ahead and anticipating responses. Some users never changed what they typed based on the actual replies. These users were actually iterating ideas by their own writing because they paid no attention to OptiMuse’s textual feedback.
@@ 5.3.3. Characteristics of Requirements in Design Practices
Although participants’ attitudes varied a lot towards OptiMuse, we observed consistency in the characteristics of requirements (requirement ambiguity, changes, and conflicts) among users using the same system. In the interview, participants agreed that the characteristics of requirements were unconsciously influenced by the difference between the two collaboration frameworks.
@@
Copilot: Strive for clarity, embrace change, and prevent conflicts. As we mentioned in Section 3, ambiguity is almost unavoidable in requirement descriptions. However, we observed how participants grapple with the challenge of being specific when using Copilot, as evidenced in their dialogues. Additionally, during interviews, they self-reported the difficulty of being as detailed as possible to ensure Copilot’s understanding. $P_{1}$ mentioned, “I was uncertain about how much detail I should provide for Copilot to ensure accurate understanding.” The requirements underwent slight changes to achieve a suitable degree or shift the focus to another aspect of the design. The background was initially instructed to be “darker,” followed by “a little bit darker,” then “still not dark enough,” and finally, “well, try to make it a little lighter.” Similar iterations of descriptions were frequently employed in dialogues between humans and Copilot. Conflicts of the requirement could not be well fixed in Copilot. First, conflicts between asynchronous requirements like “darker and brighter” were regarded as AI’s lack of ability to produce appropriate outcomes. Second, conflicts between the user and AI, namely, AI didn’t give outcomes addressing the user’s commands, were regarded as annoying mistakes. “I aimed for white titles, but for some reason, it consistently incorporated a gradient, which I found displeasing.“, $P_{10}$ stated. Conflicts within the design itself, the most tricky ones, were usually avoided in discussions. $P_{3}$ explained, “When I requested Copilot to justify both sides of all body text, the result was quite poor. I usually adjust the text box width when aligning the text. But I didn’t bother explaining this to the AI.“
@@
OptiMuse: Ambiguity sparks alternative ideas, changes and conflicts are the crucibles of creativity. There was a deliberate use of intentionally ambiguous descriptions observed in the interactions between participants and OptiMuse, as indicated by their dialogues. Furthermore, participants expressed their willingness to take OptiMuse’s opinions into account, reducing their concern about the precision of descriptions. $P_{3}$ said,“I was curious about OptiMuse’s take on this design, so I provided a vague description of my ideas and awaited its response.” OptiMuse was designed to highlight changes in requirements and conflicts, which is often where participants reported gaining novel design ideas. “I was surprised when OptiMuse questioned my commands, but its suggestions appeared reasonable. So I decided to give it a try,” $P_{8}$ added, “At least one of the selections turned out to be satisfactory. To be honest, I have never thought of doing so.” OptiMuse drove users to change requirements. In OptiMuse, users’ textual descriptions allow rapid exploration of the design space (Kulkarni et al., 2023). OptiMuse’s suggestions (transitioning from dark mode to white mode, switching theme colors, rearranging each text module based on different rules, etc) increased designers’ willingness to make bold changes. Users were prompted to decide among the changes, which encouraged reflection on each option. Users took greater tolerance towards conflicts with OptiMuse. Attitudes towards asynchronous conflicts were comparable in Copilot and OptiMuse. Nonetheless, such conflicts occurred less frequently in OptiMuse, as users were more focused on exploring new options rather than fine-tuning details. Conflicts between AI and users’ commands still happened in visual outcomes, but rather than simply dismissing them as mistakes, participants attempted to contemplate the merits of selective choices. “It felt like having an opinionated colleague. You know, negotiating and presenting options. So, I suppose there must be reasons behind whatever OptiMuse produces.”, $P_{5}$ stated. Some participants were also receptive to OptiMuse’s alternative suggestions during the dialogue.
Table 3. Expected roles of AI agents and corresponding tasks.
Roles
@@
Roles Input Interaction Output
Executor design confirm, revise optimized design
Optimizer design, requirement confirm, revise optimized design
Colleague design, requirement dialogue, select, remix selective choices
Expert design, requirement dialogue, select, remix selective choices